In the nucleus of 11Na23, the number of protons, neutrons and electrons are


        (a)    11, 12, 0   


(b)    23, 12, 11


        (c)    12, 11, 0


        (d)    23, 11, 12


Grade:10

1 Answers

Hrishant Goswami
10 years ago

(a)

Z = 11 i.e., number of protons = 11, A = 23

\      Number of neutrons = A - Z = 12

Number of electron = 0 (No electron in nucleus)

Therefore 11, 12, 0.